The circle passing through (1, -2) and touching the axis of x at (3, 0) also passes through the point (1) (2, -5) (2) (5, -2) (3) ( -2, 5) (4) ( -5, 2)

Grade:upto college level

2 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
please find the answer to your question below
let the equation of circle is (x-h)2+(y-k)2=r2.........(1)
when circle touches x axis then r=k
substituting (1,-2) in eqn(1) and r=k we get
h=3,k=-2 and r=2
(x-3)2+(y+2)2=4
From the options (5,-2) passes through the above circle
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

let the equation of circle is (x-h)^2 + (y-k)^2 = r^2.........(1)
when circle touches x axis then r=k substituting (1,-2) in eqn(1) and r=k we get h=3,k=-2 and r=2 (x-3)2+(y+2)2=4 From the options (5,-2) passes through the above circle

Thanks and Regards